OpenStudy (anonymous):

can some one check this radical expression for me to make sure i got it right?

OpenStudy (anonymous):

i got \[\frac{ 3a ^{3}}{ 2 }\]

OpenStudy (tyteen4a03):

Yes, that is correct.

OpenStudy (anonymous):

@tyteen4a03 can it be shown as \[\frac{ 3 }{2}a\] ?

OpenStudy (tyteen4a03):

@Spartan_Of_Ares No, where did the a^3 go?

OpenStudy (anonymous):

@tyteen4a03 sorry i meant \[\frac{ 3 }{ 2} a ^{3}\]

OpenStudy (tyteen4a03):

@Spartan_Of_Ares No, because the whole term 3a^3 was divided by 2, not just the 3.

OpenStudy (anonymous):

@tyteen4a03 4a^3 4a^2 \[\frac{ 3 }{2 }a ^{3}\] and \[\frac{ 3 }{2 }a ^{2}\] are the only choices

OpenStudy (tyteen4a03):

Oh sorry, went blind. 3/2 * a^3 is correct.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

2